A large stone is dropped vertically down word from a cliff overlooking the ocean if the sound of the stone hitting the water is heard 3 point40 seconds later and you know that the speed of the sound is 340 m/s how high is the cliff


1
Expert's answer
2021-03-04T15:18:25-0500

Let's first find the time that the stone takes to reach the water:


"h=\dfrac{1}{2}gt_1^2,""t_1=\sqrt{\dfrac{2h}{g}}."

Then, we can write the time that the sound of splash takes to reach the person's ear:


"t_2=\dfrac{h}{v}=\dfrac{h}{340}."

The total time can be found as follows:


"t=t_1+t_2,""\sqrt{\dfrac{2h}{g}}+\dfrac{h}{340}=3.40,""9.8h^2+253857.6h-13096092.8=0."

This quadratic equation has two roots: "h_1=25852\ m" and "h_2=51.69\ m." The correct answer is "h=51.69\ m".

Answer:

The height of the cliff is 51.69 m.